OpenStudy (anonymous):

Mary deposited $350 in a bank account that promises 2.8 percent interest compounded continuously. Approximately how many years will it take to reach a balance of $500? 1.43 years 2.80 years 5.55 years 12.77 years

OpenStudy (anonymous):

come back @saifoo.khan

OpenStudy (anonymous):

Lets try again bro

OpenStudy (anonymous):

y = years 350 x (350 x 2.8)y =500

OpenStudy (anonymous):

SORRY I meant this 350 + (350 x 2.8)y =500

OpenStudy (anonymous):

Continuous interest tells you that you will use the interest formula that involves e. A=Pert A is the future value. P is the starting value. r is the rate as a decimal. t is the number of years of the investment. You will need to use ln to get the exponent to "come down" so the equation becomes easier to manage.

OpenStudy (anonymous):

yea you right but that answer didnt tell me the years lol

OpenStudy (anonymous):

figure it out breh

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@Preetha

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

hey

OpenStudy (anonymous):

can yall help?

OpenStudy (jdoe0001):

have you covered logarithms yet?

OpenStudy (anonymous):

Thats what im on

OpenStudy (jdoe0001):

k

OpenStudy (anonymous):

lol

OpenStudy (perl):

Mary deposited $350 in a bank account that promises 2.8 percent interest compounded continuously. Approximately how many years will it take to reach a balance of $500? $$ \Large { A = Pe^{rt} \\ \therefore \\ A = 350 e^{.028t } \\~\\ \text{when does A reach 500?}\\\therefore\\ \\500 = 350 e^{.028t } } $$

OpenStudy (jdoe0001):

hmmm actauly.... lemme fix that... the term inside is a two factor

OpenStudy (jdoe0001):

but you'd end up with a log function anyhow you;'d take the log to both sides first, and then solve for "t" in the end

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I got a 5 5 5

OpenStudy (jdoe0001):

555?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Im prolly wrong

OpenStudy (jdoe0001):

\(\bf \qquad \textit{Compound Continuously Interest Amount}\\ A=Pe^{rt} \qquad \begin{cases} A=\textit{current amount}\to &500\\ P=\textit{original amount deposited}\to &\$350\\ r=rate\to 2.8\%\to \frac{2.8}{100}\to &0.028\\ t=years \end{cases} \\ \quad \\ A=Pe^{rt}\implies ln(A)=ln(Pe^{rt})\implies ln(A)=ln(P)+ln(e^{rt})\)

OpenStudy (anonymous):

so its 2.80

OpenStudy (jdoe0001):

recall the \(\bf \textit{log cancellation rule of }log_{\color{brown}{ a}}{\color{brown}{ a}}^x=x\qquad thus \\ \quad \\ ln(e^\square)\implies log_{\color{brown}{ e}}{\color{brown}{ e}}^\square \implies \square \)

OpenStudy (jdoe0001):

hehe

OpenStudy (anonymous):

maybe its 1.43

OpenStudy (anonymous):

am I wrong

OpenStudy (jdoe0001):

hmmm nope... so... solve it for "t", see what the equation gives you then put in the values given for the current balance, the principal and the rate

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

you can come back

OpenStudy (anonymous):

I got 12 years

OpenStudy (anonymous):

giving me a headache

OpenStudy (jdoe0001):

what's "t" equal to in the equation though?

OpenStudy (anonymous):

t= years

OpenStudy (anonymous):

I know t=years

OpenStudy (anonymous):

Im plugging it in

OpenStudy (anonymous):

maybe I have the wrong calculator

OpenStudy (anonymous):

500=350e^0.028(t)

OpenStudy (anonymous):

350e^0,028(12.77)

OpenStudy (anonymous):

\[350e^0.028(12.77)\]

OpenStudy (jdoe0001):

well. in "t" is not quite in "e" terms though

OpenStudy (anonymous):

so you want me to put years for e?

OpenStudy (anonymous):

this sht here

OpenStudy (jdoe0001):

hmmm nope... you're meant to solve for "t"

OpenStudy (jdoe0001):

and then you can just replace the values for "r" and "P" and "A" for the given amounts for rate, principal and amount,, respectively

OpenStudy (anonymous):

t = (1/r)(A/P - 1)

OpenStudy (anonymous):

(1/0.028)(500/350-1)

OpenStudy (anonymous):

I dont know Im done

OpenStudy (anonymous):

well it is 12.77